### v1

@@ 4.1.2 An elliptical credible set for $\theta$
An optimal method to create credible set for $\theta$ is the highest posterior density (HPD) method. In this method, any point $\theta_{*}$ in the parameter space of $\theta$ is considered for possible inclusion in the credible set $\Theta_{H}$ if the posterior density $\pi_{M}(\theta_{*}|y_{1},\cdots,y_{m})$ is higher than the density at any point outside the set $\Theta_{H}$. We continue adding such $\theta_{*}$ values to the HPD set until the credible level of the set is at least $(1-\alpha)$. The method is intuitively sound since it includes the more likely $\theta$ values in terms of the posterior distribution. Obviously, an HPD credible set is of the form $\{\theta:\pi_{M}(\theta|y_{1},\cdots,y_{m})\geq t\}$, where the positive threshold $t$ depends on $y,\alpha$, and is chosen by satisfying the credible level $(1-\alpha)$. In terms of “volume”, among all $(1-\alpha)$ credible sets, the HPD credible set has the lowest volume. For the unstructured Bayesian model in Section 2.1, an HPD credible region is of the form
$\left\{\theta:\sum_{i=1}^{m}(y_{i}-\theta_{i})^{2}/D_{i}\leq c\right\}$ (9)
@@
for a suitable constant $c$. The shape of this credible set is elliptical. The defining equation of the HPD credible set above shows that if a point $\theta$ is inside the set, the components of that point cannot vary freely since a larger value of $(y_{i}-\theta_{i})^{2}$ will make a larger value of $(y_{j}-\theta_{j})^{2}$ for $j\neq i$ less likely. This is in direct contrast to the Cartesian product credible region in subsection 4.1.1. In the frequentist approach, Klein et al. (2020) used a confidence region which is identical under the unstructured model to the credible set determined by Cartesian product. This confidence set allows to line up the individual intervals to determine overall ranking of the means. It should be mentioned that for the frequentist approach it is possible to get an elliptical confidence set of the form in (9) by inverting an appropriate likelihood-based acceptance region. However, due to interdependence of values of the components of $\theta$ that are allowed in the region, it would not be possible to line up these components to determine their relative rankings. This is a major hurdle with the use of the optimal likelihood-based confidence region to determine confidence region for overall ranking $\check{r}$. However, this difference in shape between the rectangular credible region determined by the Cartesian product and the “elliptical” HPD region does not pose any challenge to the ranking problem in our Bayesian setup. Since we can deal with the values of $\theta$ in the credible set directly, determining the overall ranking of the components is not a problem. This presents a major advantage to the Bayesian approach. The credible set in equation (9) depends on the Mahalanobis distance (MD) function
@@
where $y=(y_{1},\cdots,y_{m})$ is the posterior mean vector and $D=\mbox{diag}(D_{1},\cdots,D_{m})$ is the posterior variance-covariance matrix of $\theta$. Based on a posterior sample ${\cal F}_{UB}$ of $\theta$, we can compute the set of MD values $\{M(\theta,y,D):\theta\in{\cal F}_{UB}\}$ and take the $(1-\alpha)$ quantile of these values for the cut-off $c$, and construct an empirical HPD credible set given by
@@
For the more general HB model in Section 2.2, which is completed by the noninformative improper prior $\pi(\beta,A)$ in equation (6), for large $m$ the posterior pdf $\pi_{HB}(\theta|y)$ is approximately multivariate normal (it follows heuristically by noting that the conditional posterior distribution of $\theta$ given $A$ is a multivariate normal distribution, and applying the Laplace approximation based on asymptotic normality of the marginal posterior distribution of $A$). Consequently, an HPD credible set for $\theta$ will be approximately elliptical in shape, which can be defined by the Mahalanobis distance function based on the posterior mean vector and posterior variance-covariance matrix of $\theta$. Specifically, we can construct empirically an HPD credible set based on the sample of $\theta$ values, ${\cal F}_{HB}$, from the posterior pdf. To this end we compute estimates of the posterior mean and the variance-covariance matrix, given by
@@
respectively. Next, we compute the Mahalanobis distances $d^{(s)}_{HB}=M(\theta^{(s)},\hat{\theta}_{HB},V_{HB})$ for $s=1,\cdots,S$. Again, if $c$ is the $(1-\alpha)$-quantile of these distances, we can create an empirical credible set ${\cal F}_{HB,\alpha,y}$for $\theta$, similar to ${\cal F}_{UB,\alpha,y}$, by using $\hat{\theta}_{HB}$ and $V_{HB}$ in place of $y$ and $D$, respectively. This credible set is a sample-based approximation to the HPD credible set for $\theta$ with credible coefficient $(1-\alpha)$. We remark that it is not surprising that the observed cut-off point $c$ is approximately equal to the $(1-\alpha)$-quantile of a chi-square random variable with $m$ degrees of freedom.
5 Joint distribution of overall ranking on a $(1-\alpha)$ credible set
We have underscored it before the importance of accounting for uncertainty for a set of released ranks based on some point estimates of the subpopulation means. Klein et al. (2020) addressed this problem in their fundamental paper from a frequentist perspective. We now present a simple and straightforward solution to this problem by using a Bayesian approach. In Bayesian statistics, inference for any function, possibly vector-valued, of the parameter can be conducted by computing a multi-dimensional relative frequency distribution of the function computed based on a sampling-based approach by drawing a reasonably large random sample of $\theta$ from its posterior distribution. To implement our approach, for a given $\alpha$, we take some large $S$ and generate a sample $\theta^{(1)},\cdots,\theta^{(S)},$ from the posterior distribution. From this sample, we select $\theta^{(s_{1})},\cdots,\theta^{(s_{K})}$, which we denoted earlier as ${\cal F}_{M,\alpha,y}$, based on procedures explained above, where $K$ is the integer nearest to $S(1-\alpha)$. For each sample in ${\cal F}_{M,\alpha,y}$, we then create a two-way $m\times m$ table, columns as the populations or subjects to be ranked and rows as the ranks of the components of $\theta$, explained now. For each sample such as $\theta^{(s_{1})}$, we start with an $m\times m$ null matrix. If the $i$th component of $\theta^{(s_{1})}$ does not tie with any other component and has rank $j_{i}$, we put a $1$ in the $j_{i}$th row of the $i$th column. If two components $k$ and $l$ tie for ranks $j$ and $j+1$, we replace elements $(k,j),(k,j+1),(l,j),(l,j+1)$ by $1/2$. We proceed similarly with ties among more than two components. In this way for all $K$ elements in ${\cal F}_{M,\alpha,y}$, we complete $K$ tables, and we take a weighted average of these tables, where the weights sum to $K$ so that the resulting table has its each row and column sum to $K$. Two possible choices of the weights are (a) equal weights, or (b) proportional to $\exp[-M(\theta^{(s)},\cdot,\cdot)/2]$, where $M(\theta^{(s)},\cdot,\cdot)$ is the suitable Mahalanobis distance. By converting these entries of this table to fractions so that all rows and all columns sum to 1, we obtain a double stochastic square matrix. The $k$th row assigns probabilities to various subpopulations or players to hold the rank $k$. The $i$th column of this matrix is providing an empirical posterior distribution of rank ${\check{r}}_{i}$, the rank of the $i$th population, player or object. We name this probability distribution as the credible distribution. For future reference we write this probability distribution of ${\check{r}}_{i}$ by $\pi_{{\check{r}}_{i},M,T,y}$. We added an additional subscript “T” to indicate the type of credible sets: for Cartesian T=C, and for elliptical T=E. With respect to this distribution, for a function $g({\check{r}}_{i})$, we use $E^{\pi_{{\check{r}}_{i},M,T,y}}[g({\check{r}}_{i})]$ to denote its posterior expectation.
6 Comparison of the set estimators based on size measures
@@
Two types of set estimators for the mean vector $\theta$ are playing the critical part in accounting for uncertainty in the inference for overall ranking of the component means. To this end, in their frequentist approach Klein et al. (2020) used a confidence region for the mean vector which is the Cartesian product of appropriate confidence intervals of the component means. In an effort to construct a Bayesian solution which closely resembles this frequentist solution, we have also developed a credible set, an orthotope or a hyperrectangle, which is a Cartesian product of suitable credible intervals for the individual means. As an alternative to this orthotope shaped credible set, where each component of $\theta$ varies freely, we have also proposed HPD sets for various Bayesian models that used here. It is intuitively clear that HPD sets would be more “compact” than sets of other shapes in terms of reasonable size measures. In this section, we motivate and compute two size measures of various set estimators that have comparable confidence or credibility level, and compare the sets based on these measures. Two size measures that appear to be appropriate for set estimators are their Lebesgue volumes and average lengths of the “sides” determining the set. Here, a set estimator $C$ of the Cartesian form is given by $[L_{1},U_{1}]\times\cdots\times[L_{m},U_{m}]$. The volume, $V_{C}$, and the average length $L_{C}$, of this set, respectively, are
@@
$E=\{\theta:(\theta-\theta^{*})^{T}K(\theta-\theta^{*})\leq c\},$ (12)
where $K$ is an $m\times m$, appropriate p.d. matrix, $\theta^{*}$ is an appropriate center, and $c>0$ is a suitable cut-off, all known. The Lebesgue volume $V_{E}(K,m,c)$ of the set $E$ has been provided in equation (16) in the Appendix. A suitable length measure of an elliptical region is not available. We have derived in equation (19) in the Appendix a reasonable length measure of an elliptical region. From that consideration an average length of the set $E$ is
@@
where the expression of $L_{M,i}$ is provided by equation (19) in the Appendix.
@@
For HPD or approximate HPD credible sets, the matrix $K$ in equation (12) corresponds to $V_{HB}^{-1}$ in HB models and $K=D^{-1}$ in the unstructured Bayesian model. The matrices $V_{HB}$ and $D$ are the posterior variance-covariance matrices for the two models, respectively.
7 Ranking of baseball players: An example by Efron and Morris
Efron and Morris (1975) considered batting averages of 18 major league baseball batters from their first 45 at bats in the 1970 baseball season to predict their performances in the remainder of that season. Let $\theta_{i}$ be the unknown true batting average of the $i$th player during that season. A reasonable representative value for $\theta_{i}$ was the player’s average ($\omega_{i}$) in the remainder of the season after his first 45 at bats. The sample proportion of hits $Y_{i}$ from first 45 at bats is approximately normal with mean $\theta_{i}$ and variance $D_{i}$. In our analysis, we use the estimated variance, given by $D_{i}=Y_{i}(1-Y_{i})/45.$ Treating the available $\omega_{i}$ as the “gold standard” value of $\theta_{i}$, for the direct estimates $y_{i}$’s the total empirical squared error (TESE), $\sum_{i=1}^{18}(y_{i}-\omega_{i})^{2}$, is $0.07537$, and the same error for the HB and the EBLUP estimates of $\theta_{i}$’s are $0.02368$ and $0.02356$, respectively. The HB estimates have $3.18$ times smaller TESE than the direct estimates, and the EBLUP estimates have $3.20$ times smaller. Efron and Morris (1975) used the arc-sin transformed parameters $\psi_{i}=arcsin(2\theta_{i}-1)$, for the transformed data $\hat{\psi}_{i}=arcsin(2Y_{i}-1)$ in their EB analysis. This transformation resulted in equal approximate sampling variances, $1/45$, for all 18 $\hat{\psi}_{i}$’s. They applied Stein’s shrinkage estimation method based on an exchangeable model to calculate ${\hat{\psi}}_{J-S,i}$, the James-Stein estimate of $\psi_{i}$. Treating the players’ batting averages $\omega_{i}$ for the remainder of the season are closely related to true batting averages $\theta_{i}$’s, they calculated the TESE of prediction for the direct estimates $\hat{\psi}_{i}$ from $arcsin(2\omega_{i}-1)$ is 0.3902, which is about 3.5 times of 0.1113, the TESE of prediction for the James-Stein estimates ${\hat{\psi}}_{J-S,i}$. This application showed a substantial efficiency of Stein’s method. The efficiency for HB estimators of $\psi_{i}$ based on the same model is about $3.36$, which is comparable. Note that these numbers are comparable to $3.18$ and $3.20$, we obtained above for EBLUP and HB predictors of $\theta_{i}$’s without transformation.

### v2

@@ 4.1.2 An elliptical empirical credible set for $\theta$
An optimal method to create credible set for $\theta$ is the highest posterior density (HPD) method. A point $\theta_{*}$ in the parameter space of $\theta$ is eligible for inclusion in an HPD credible set $\Theta_{H}$ if the posterior density $\pi_{M}(\theta_{*}|y_{1},\cdots,y_{m})$ is no less than the density at any point outside the set $\Theta_{H}$. We continue adding such $\theta_{*}$ values to the HPD set until the credible level of $\Theta_{H}$ is at least $(1-\alpha)$. The method is intuitively sound since it includes the more likely $\theta$ values in terms of the posterior distribution. Obviously, an HPD credible set is of the form $\{\theta:\pi_{M}(\theta|y_{1},\cdots,y_{m})\geq t\}$, where the positive threshold $t$ depends on $y,\alpha$, and is chosen by satisfying the credible level $(1-\alpha)$. In terms of Lebesgue volume, among all $(1-\alpha)$ credible sets, the HPD credible set has the lowest volume.
For the UB model in Section 2.1, for a constant $c$ an HPD credible set is of the form
$\left\{\theta:\sum_{i=1}^{m}(y_{i}-\theta_{i})^{2}/D_{i}\leq c\right\}.$ (9)
@@
The HPD credible set above, elliptic shape, shows that if a point $\theta$ is inside the set, the components of that point cannot vary freely since a larger value of $(y_{i}-\theta_{i})^{2}$ will make a larger value of $(y_{j}-\theta_{j})^{2}$ for $j\neq i$ less likely. This is in direct contrast to the Cartesian credible region in Subsection 4.1.1. It should be mentioned that for the frequentist approach it is possible to get an elliptical confidence set of the form in (9) by inverting an appropriate likelihood-based acceptance region. However, due to interdependence of values of the components of $\theta$ that are allowed in the region, it would not be possible to line up these components to determine their relative rankings. This is a major hurdle with the use of the optimal likelihood-based confidence region to determine confidence region for overall ranking $\check{r}$. However, shape of a credible region poses no challenge to the ranking problem in our Bayesian setup. Since we can deal with the values of $\theta$ inside a credible set directly, determining the overall ranking of the components is not a problem. It is a major advantage to the Bayesian approach.
The credible set in equation (9) depends on the Mahalanobis distance (MD) function
@@
where, for the UB model, $y$ is the posterior mean and $D=\mbox{diag}(D_{1},\cdots,D_{m})$ is the posterior variance-covariance matrix of $\theta$. Based on a posterior sample ${\cal F}_{UB}$ of $\theta$, we can compute the set of MD values $\{M(\theta,y,D):\theta\in{\cal F}_{UB}\}$ and take the $(1-\alpha)$ quantile of these values for the cut-off $c$, and construct an empirical HPD credible set given by
@@
For the more general HB model in Section 2.2, for large $m$ the posterior pdf $\pi_{HB}(\theta|y)$ is approximately multivariate normal (it follows heuristically by noting that the conditional posterior distribution of $\theta$ given $A$ is a multivariate normal distribution, and applying the Laplace approximation and consistency of the posterior mode of $A$). Thus, an HPD credible set for $\theta$ will be approximately elliptic, defined by the Mahalanobis distance function based on the posterior mean and variance-covariance of $\theta$. Specifically, we can construct an empirical HPD credible set using the sample of $\theta$ values, ${\cal F}_{HB}$, from the posterior pdf. To this end we compute estimated posterior mean and variance-covariance matrix, given by
@@
respectively. Next, we compute the Mahalanobis distances $d^{(s)}_{HB}=M(\theta^{(s)},\hat{\theta}_{HB},V_{HB}),$ $s=1,\cdots,S$. With $c=(1-\alpha)$-quantile of these distances, we create an empirical credible set ${\cal F}_{HB,\alpha,y}$ for $\theta$, similar to ${\cal F}_{UB,\alpha,y}$, based on $d^{(s)}_{HB}$ values. This set is a sample-based empirical representation of the HPD credible set for $\theta$ with credible coefficient $(1-\alpha)$. Note that $c\approx\chi^{2}(1-\alpha,m)$, the $(1-\alpha)$-quantile of chi-square distribution with $m$ degrees of freedom.
5 Joint distribution of overall ranking on a $(1-\alpha)$ credible set: credible distributions
We have underscored the importance of accounting for uncertainty for a set of released ranks based on some point estimates of the subpopulation means. We now present a simple and straightforward solution using a Bayesian approach. Bayesian inference for any function, possibly vector-valued, of the parameter can be conducted by computing a multi-dimensional relative frequency distribution of the function based on a reasonably large sample of $\theta$ from its posterior distribution.
From a sample $\theta^{(1)},\cdots,\theta^{(S)}$ from $\pi_{M}(\theta|y)$ we select $\theta^{(s_{1})},\cdots,\theta^{(s_{K})}$, to be denoted by ${\cal F}_{M,\alpha,y}$. For each sample in ${\cal F}_{M,\alpha,y}$, we then create an $m\times m$ two-way table, columns as the entities to be ranked and rows as the ranks of the entities. For the sample $\theta^{(s_{1})}$, we start with an $m\times m$ null matrix. If the $i$th component of $\theta^{(s_{1})}$ does not tie with any other component and has rank $j_{i}$, we put a $1$ in the $j_{i}$th row of the $i$th column. If two components $k$ and $l$ tie for ranks $j$ and $j+1$, we replace elements $(k,j),(k,j+1),(l,j),(l,j+1)$ by $1/2$. We proceed similarly with ties among more than two components. In this way for all $K$ values of $\theta$ in ${\cal F}_{M,\alpha,y}$, we complete $K$ tables, and we take a weighted average of these tables, where the weights sum to $K$ so that the resulting table has its each row and column sum to $K$. Two possible choices of the weights are: (a) equal weights, or (b) proportional to $\exp[-M(\theta^{(s)},\cdot,\cdot)/2]$, where $M(\theta^{(s)},\cdot,\cdot)$ is the suitable Mahalanobis distance. Throughout this paper, the “unweighted” method actually corresponds to equal weights (case (a) above).
By converting the entries of this table to fractions so that all rows and all columns sum to 1, we obtain a double stochastic square matrix, which plays a pivotal role in this paper. Its $k$th row assigns probabilities to various entities to hold the rank $k$. The $i$th column of this matrix is providing an empirical posterior distribution of rank ${\check{r}}_{i}$, the rank of the $i$th entity. We name this probability distribution as the credible distribution (CD). For later use we write this probability distribution of ${\check{r}}_{i}$ by $\pi_{{\check{r}}_{i},M,T,y}$. We added the subscript “T” to indicate the type of credible sets: T=C for Cartesian, and T=E for elliptical. With respect to this distribution, for a function $g({\check{r}}_{i})$, we use $E^{\pi_{{\check{r}}_{i},M,T,y}}[g({\check{r}}_{i})]$ to denote its posterior expectation.
6 Comparison of set estimators of $\theta$ by size
@@
Two types of set estimators for the mean vector $\theta$ are playing the critical part in accounting for uncertainty in the inference for overall ranking. Klein et al. (2020) used a confidence region for the mean vector which is the Cartesian product of appropriate confidence intervals of the component means. In our Bayesian effort to closely resemble the frequentist set for $\theta$, we have also developed a credible set, which is an orthotope or a hyperrectangle, a Cartesian product of suitable credible intervals for the individual means. As an alternative to this Cartesian set, where each $\theta_{i}$ varies freely, we have also proposed HPD sets for $\theta$. It is intuitively clear that in terms of reasonable “size” measures HPD sets would be more “compact” than sets of other shapes. In this section, we use two obvious measures of set estimators of $\theta$ with comparable confidence or credibility level, and compare various sets.
Two size measures that appear to be appropriate for a set estimator is its Lebesgue volume and average length of the “sides” determining the set. A Cartesian set estimator $C$ is given by $\prod_{i=1}^{m}[L_{i},U_{i}]$. Its volume, $V_{C}$, and average length, $L_{C}$, respectively, are
@@
$E=\{\theta:(\theta-\theta^{*})^{T}\Phi(\theta-\theta^{*})\leq c\},$ (12)
where $\theta^{*}$ is an appropriate center, $\Phi$ is an appropriate $m\times m$, p.d. matrix, and $c>0$ is a suitable cut-off, all known. The Lebesgue volume $V_{E}(\Phi,m,c)$ of the set $E$ is well known, and its expression is shown in equation (S.2) in the Appendix. A suitable length measure of an elliptical region is not available. We have attempted a reasonable length measure in the Appendix. The length measure, denoted $L_{E}$, is
@@
where the “length” $L_{M,i}$ is provided by equation (S.5) in the Appendix.
@@
For HPD or approximate HPD credible sets, the matrix $\Phi$ in equation (12) corresponds to $V_{HB}^{-1}$ in HB models and $\Phi=D^{-1}$ in the UB model.
7 Comparison of frequentist and Bayes solutions to the rank inference problem
Frequentist and Bayes solutions to the set inference problem for ranks are very different. The frequentist KWW solution is a Cartesian product of subsets of $\{1,2,\cdots,m\}$. The Bayesian solution is through probability distributions, the CDs, on $\{1,2,\cdots,m\}$. Given their different nature, these solutions can be compared only indirectly.
7.1 Comparison via entropy
A KWW confidence interval of $\check{r}_{i}$, rank of entity $i$, is an interval $\{a_{i}+1,\cdots,a_{i}+n_{i}\}$ for a positive integer $n_{i}$ and a nonnegative integer $a_{i}$. A CD of $\check{r}_{i}$ is a probability distribution on $1,\cdots,m$. These solutions are not directly comparable. In the absence of any other information for a KWW set, we need to assume the possible values of $\check{r}_{i}$ are uniformly distributed on the KWW set, and calculate entropy or other deviation measure that would compare the “concentrations” provided by these solutions. The entropy of a probability distribution measures its uncertainty or concentration. The entropy of a discrete distribution $f$ is
$E=-\sum_{x:f(x)>0}f(x)\log(f(x)),$
see, for example, p.91, Berger (1985). Lower entropy means a distribution is more concentrated or informative. The lowest entropy is 00, attained only by a degenerate distribution. We apply this concept to our applications.
7.2 Comparison via cardinality
To evaluate a deviation measure mentioned above, we can proceed as follows. Again, for a discrete random variable $W$ taking values $a_{i}+1,\cdots,a_{i}+n_{i}$, each with probability $1/n_{i}$, it easily follows that $4E\Big{(}|W-\mu|\Big{)}=n_{i}-n_{i}^{-1}C$, where $C$ is zero if $n_{i}$ is even, and $C=1$ if $n_{i}$ is odd, where $\mu=E(W)$. An importance of this result is that if the KWW rank set for $\check{r}_{i}$ contains $n_{i}$ points $\{a_{i}+1,\cdots,a_{i}+n_{i}\}$, then $n_{i}\approx 4E|W-\mu|$.
8 Ranking of the US states based on commuting times
In a pioneering article that motivated our study, Klein et al. (2020) applied their frequentist approach to rank fifty states of the U.S. and DC by mean commuting times of workers sixteen or older and not working from home. They used survey data collected from the ACS. In Figure 1, we recreated Figure 1 of Klein et al. (2020) that depicted the frequentist solution of the confidence region for ranking from equation (8). Possible ranks for state $i$ in this set are shown by a pink line segment stretching from $|\Lambda_{Li}|+1$ to $|\Lambda_{Li}|+|\Lambda_{Oi}|+1$. From the figure, for example, the possible ranks for the state ID are $4-9$. On the other hand, the states which can hold the rank 9 are the states WY, AK, MT, IA, KS and ID (the pink line segments for these states intersect the horizontal line for Ranks = 9). The rank based on direct estimate $y_{i}$ is shown by a black dot at its rank $r$. If two states tie for ranks $r,r+1$, their tied ranks are shown by a black line segment joining these states at rank $r+1$. See, for example, SD and ND direct estimates tie for ranks 1 and 2.